OpenStudy (wintersuntime):

two car rental agencies are competing for business. Mertz charges $35 and $0.15 per mile, and mavis charges $20 and $0.45 per mile. Determine when the two agencies will charge the same amount to rent a car. How much will they charge ?

OpenStudy (wintersuntime):

Using any method

OpenStudy (nubeer):

Let mile = m For mertz it would charge = 35 + 0.15*m can u find what mavis would charge ?

OpenStudy (wintersuntime):

m= 20+ 0.45m ??

OpenStudy (wintersuntime):

@nubeer

OpenStudy (nubeer):

yes.. now put both equations equal to each other 35+0.15m = 20+0.45m now solve it